OpenStudy (anonymous):

limz=4z2−64−−−−−−√÷z−4

OpenStudy (mathteacher1729):

I don't know what you're asking?

OpenStudy (anonymous):

that is square root

OpenStudy (anonymous):

This is not a maths problem at all. It's just spam.

OpenStudy (mathteacher1729):

\[\lim_{z\to??} \frac{4z^2-64}{\sqrt{z-4}}\] Is that what you're asking?

OpenStudy (anonymous):

no overall square root

OpenStudy (anonymous):

z -> 4

OpenStudy (mathteacher1729):

I'm giving you the benefit of the doubt here, but I have no idea what you're asking.

OpenStudy (sriram):

the answer is tending to 0

OpenStudy (mathteacher1729):

Oh, I think I see now: \[ \lim_{z\to 4} \sqrt{\frac{4z^2-64}{z-4}} \]

OpenStudy (anonymous):

2 sqrt (z+4), z to 4 is 4 sqrt 2.

OpenStudy (sriram):

i guess estudier is correct

OpenStudy (mathteacher1729):

People. Please. Use the "Equation" button on the lower-left of the text box.

OpenStudy (mathteacher1729):

\(\lim_{z\to 4} \sqrt{\frac{4z^2-64}{z-4}}=\lim_{z\to 4} 2\sqrt{\frac{z^2-16}{z-4}}=\lim_{z\to 4}2\sqrt{\frac{(x-4)(x+4)}{z-4}}\lim_{z\to 4}2\sqrt{x+4}\)

OpenStudy (mathteacher1729):

\[\begin{align}\lim_{z\to 4} \sqrt{\frac{4z^2-64}{(z-4)}}=&\lim_{z\to 4} 2\sqrt{\frac{z^2-16}{(z-4)}}\\=&\lim_{z\to 4}2\sqrt{\frac{(z-4)(z+4)}{(z-4)}}\\=&\lim_{z\to 4}2\sqrt{z+4}\\=&2\sqrt{4+4}\\=&2\sqrt{8}\\=&4\sqrt{2}\end{align}\] DONE. :) (see attached)

OpenStudy (anonymous):

Pain, isn't it?